A piston - cylinder device contains 0.15 kg of steam at 1 MPa and 320 degree C. The steam now expand to a final state of 150 kPa and 160 degree C. Energy losses via heat transfer from system to surroundings are estimated to be 3KJ, while the contact temperature with the surroundings is 25 degree C.Assuming: T0=25c and P0=100kPa,Determine:

A: the available energy of the steam at the initial and final states

B: the work done by the system

C: THE exergy (available energy) transfer in the work and heat interaction.

D: THE exergy destruction (irreversibility) for the process
Answer #1

State 1(us?ng steam tables) u230 KT(E KT v,0. 26g P MPa h 3010 T320c State 2 2600 kg 3 2 m k P = 150 K Pa h Dead state = \০4m(h-h) b) hork done by the sys tem W o.5(8090- 2190) c) Exergy tranfer, ax= X2- X7 = +6.56 -I67- 3 5 -3 o.79 kJ d) energy bal